# Modeling the One-to-Many Property in Open-Domain Dialogue with LLMs

### v1

@@ 1 Introduction
Prior research has primarily focused on modeling the o2m property using probabilistic learning frameworks, such as the Conditional Variational Auto-Encoder (CVAE) Shen et al. (2017); Zhao et al. (2017), to enhance response diverity. These methods typically condition the response on both the dialogue context and a randomly sampled latent variable, capturing the variability in conversational responses and effectively modeling the o2m property. Other approaches include randomized architectures Lee et al. (2022b), Wasserstein Auto-encoders Gu et al. (2018), and Bayesian architectures Lee et al. (2023). These studies illustrate that while explicitly modeling the o2m property of OD significantly enhances response diversity, there is typically a trade-off with contextual coherence Sun et al. (2021); Lee et al. (2022a).
Recent advancements in Large Language Models (LLMs) have made it increasingly impractical to model the o2m property using probabilistic approaches, primarily due to the immense scale of modern LLMs Zhao et al. (2023). These frameworks typically employ a pretrained LLM as the decoder, which is fine-tuned along with additional network components responsible for generating the latent distribution. This process becomes highly resource-intensive given the scale of these LLMs. Moreover, many state-of-the-art LLMs operate as black boxes with undisclosed parameters. Therefore, in the context of LLMs, adopting probabilistic frameworks for generating responses to model the o2m property has become largely impractical.
@@
To facilitate MRG, we curate o2mDial, a novel conversational dataset that explicitly captures the o2m property of OD. To create o2mDial, we leverage the DailyDialog corpus. First, we sample 500 dialogues (three to six turns) from the training set of the DailyDialog corpus. In this paper, for MRG, we fix $n=5$. In other words, we aim to generate a set of five lexically and semantically distinct, yet contextually coherent responses. Unlike prior datasets that feature multiple reference responses Hedayatnia et al. (2022); Sai et al. (2020); Gupta et al. (2019) that rely on the same LLM to generate every reference response, we use five distinct LLMs to simulate five different agents, with each LLM generating one response. As far as possible, this ensures the semantic and lexical uniqueness of each response. Based on our resource constraints, we selected the following five LLMs: 1)gpt-3.5-turbo OpenAI (2021); 2)llama2-70b-chat et al. (2023); 3)mixtral-8x22b et al. (2024); 4) StableVicuna13b Chiang et al. (2023); 5) Flan-T5-xxl Chung et al. (2022). Additionally, to construct a separate test set for MRG evaluation, we sample another 100 dialogue samples from the test of the DailyDialog corpus. Similar to the training set, each turn consists of a dialogue context (three to six turns), and a set of five distinct and contextually coherent responses.
@@
{tblr}
column2 = c, hline1,4 = -, # samples & 500(train)/100(test)
@@ 3 Multi-Response Generation
It should be noted that even if all $n$ responses convey similar ideas, they can still be lexically unique by using different vocabularies or structures. Thus, a set of responses can exhibit high inter-response lexical diversity while maintaining low inter-response semantic diversity. For example, the statements "A heavy downpour is happening" and "There’s a strong rainfall occurring" are lexically unique due to their different words and structures. However, the semantic content—that it is raining heavily—remains the same, indicating high semantic similarity. Our goal is to generate responses with both high inter-response lexical diversity and high inter-response semantic diversity, while ensuring contextual coherence. Prior work in MRG involve utilizing different sampling strategies, which produce responses with low semantic diversity, and pre-specified dialogue acts, which are significantly more complex to implement Sakaeda and Kawahara (2022).
@@ 3.1 Methods
Chain-of-thought (CoT) Prompt Chain-of-Thought (CoT) prompting Wu et al. (2023) involves prompting the model to generate intermediate steps or explanations in addition to the final answer. In our case, we prompt the LLM to explain how each response differs from the other responses. We hypothesize that by prompting the model to identify the differences between generated responses, the model would be more inclined to generate lexically and semantically diverse responses. The prompt template is provided in Figure 4, located in the Appendix.
Prompt Chaining (PC) Prompt Chaining (PC) Sun et al. (2024) typically involves dividing a task into smaller subtasks and executing them sequentially using prompts, where the output of one prompt serves as the input for the next. In our approach, we use PC to guide the LLM in generating a set of $n$ unique responses one by one. The process begins with an initial prompt $\mathbf{P}_{0}$ that asks the LLM to generate a response to the dialogue context. Subsequent prompts ($\mathbf{P}_{1}\cdots\mathbf{P}_{n-1}$) instructs the LLM to generate contextually coherent responses that differ semantically and lexically from every response generated by the previous prompts, which are included in the current prompt as input. We hypothesize that by decomposing the task of MRG into $n$ smaller subtasks, the LLM can more effectively ensure both lexical and semantic uniqueness across the responses. However, it is important to note that PC requires multiple inferences from the LLM. Therefore, generating $n$ responses requires $n$ separate inferences, which could impact the feasibility and efficiency of this approach in the real world. The prompt template is provided in Figure 5, located in the Appendix.
@@
Instruction Tuning (IT) In addition, we also conduct IT via QLoRADettmers et al. (2023) using the collected corpus. IT with QLoRA was performed using a batch size of 32, a learning rate of 2e-4, 4 epochs, a rank of 16, an alpha of 32, and a dropout of 0.05. The instruction used for IT is identical to the zero-shot variant (prompt consists of only the instruction without any demonstration examples) of the FS prompt (Figure 3).
@@ 3.2 Evaluation
Inter-response Semantic & Lexical Diversity In the context of open-domain dialogue, response diversity is typically measured via the Distinct metric, which is typically calculated by dividing the number of unique n-grams by the total number of n-grams. However, in our case, we aim to quantify the relative diversity of a set of $n$ responses. In other words, we would like to measure, on average, how different each response is from the other $n-1$ responses. Additionally, based on our definition, it would be ideal if semantic and lexical diversity can be evaluated separately. To this end, we define two separate metrics each accounting for either inter-response semantic or lexical diversity respectively: the inter-response semantic diversity score ($d_{sem}(R_{n})$) and inter-response lexical diversity score ($d_{lex}(R_{n})$).
@@
Contextual Coherence For our task, the overall contextual coherence of a set of responses can be attained by averaging the individual scores attained by each of the $n$ responses in $R_{n}$ (Algorithm 3). We employ two contextual coherence metrics: the Utterance Entailment (UE) score Lee et al. (2022a), and the UniEval-dialog coherence score Zhong et al. (2022). The UE score involves framing the task of contextual coherence evaluation as a Natural Language Inference (NLI) task. For each utterance in the dialogue context and the corresponding generated response, a NLI model assesses whether the response entails, contradicts, or is neutral with respect to the utterance. For each response $r_{i}$ from $R_{n}$, $UE(r_{i})=\frac{1}{m}\sum_{j\in m}NLI(r_{i},d_{j})$. The UE score of a set of $n$ responses is a continuous number between 0 and 1, where a greater value would indicate greater contextual coherence. The UniEval-dialog is a LLM-based approach which involve re-framing response evaluation as a boolean question and answer task. Essentially a LLM is finetuned and prompted to generate either ’Yes’ or ’No’ to the question: ’Is this a coherent response given the dialogue history?’. Hence, for the UniEval-dialogue coherence metric, each response is assigned a score of 1 if ’Yes’ is generated or 0 if ’No’ is generated.
However, evaluating the contextual coherence of OD dialogue responses remains a challenging problem and an active area of research due to the o2m property Li et al. (2016). Hence, in our experiments, we conduct a human evaluation to further support our findings.
@@ 4 Preference-based Selection (PS)
PS involves selecting the final response $r_{f}$ from $R_{n}$ based on human preference. Unlike traditional open-domain dialogue criteria such as coherence, diversity, engagingness, naturalness, or fluency, human preference covers broader factors like helpfulness, harmlessness, and interestingness Li et al. (2024). We prioritize human preference for three key reasons. Firstly, MRG already ensures coherence and diversity within $R_{n}$. Secondly, existing metrics fall short in capturing the full complexity of human preferences, as they address only specific aspects of response quality Jiang et al. (2024). Thirdly, modern LLMs are largely capable of generating fluent, natural, and engaging responses Zhao et al. (2023).
However, in addition to human preference, the contextual coherence of the response should still be considered during selection. Hence, for PS, we aim to design an Open-domain Dialogue Response Preference (ODRP) model that assigns a scalar score to each response in $R_{n}$ based on human preference. To achieve this, we leverage an open-source preference model from OpenAssistant HuggingFace (2021) based on deberta-v3-large commonly used for Reinforcement Learning with Human Feedback (RLHF) et al. (2022). Such models are typically trained on preference datasets derived from tasks such as summarization et al. (2020) and question answering et al. (2021), or curated specifically to prevent harmful behavior et al. (2022). Hence, to fine-tune the preference model for open-domain dialogue, we construct a new preference dataset from the corpus described earlier.
Preference datasets consist of comparisons between two responses given the same prompt (dialogue context $D$ in our case). Extending o2mDial, we construct a preference dataset for fine-tuning by engaging annotators to label the preferred response $y_{c}$ and the rejected response $y_{r}$ (based on which they would prefer from a conversation partner) for every possible pair of responses from $R_{n}$, resulting in ($n\choose 2$ pairs per set. As per Ouyang et al. (2022), we consider every pair from $R_{n}$ as a single batch. The preference model is then fine-tuned on the following contrastive loss function:
@@
We then proceed to fine-tune the preference model via QLoRA Dettmers et al. (2023) for two epochs with AdamW (lr=2e-4). After MRG, the ODRP model assigns a score to each response in $R_{n}$, and $r_{f}$ is the response with the highest score:
@@
Additionally, we introduce a variant of the ODRP model finetuned on a subset of the corpus selected via hard negative sampling Robinson et al. (2021). Specifically, we apply the base preference model to the dataset and deliberately extracted samples ( 50%) on which the base model performed the worst (assigned a similar score for both $y_{c}$ and $y_{r}$ or assigned a higher score to $y_{r}$). We finetuned this variant of the ODRP model for four epochs instead.
@@ 5.1 Corpora
For evaluation, we use two main datasets: DailyDialog Li et al. (2017) and EmpatheticDialogs Rashkin et al. (2019). DailyDialog features diverse, open-domain multi-turn conversations, while EmpatheticDialogs focuses on responses to emotionally grounded events. In our experiments, the dialogue agent’s task is to generate responses based solely on the context of the ongoing conversation. We do not use any additional information such as response labels (e.g., emotion, topic, or style) or speaker labels.
@@ 5.2 Implementation
We generate five responses per context ($n=5$) using TinyLlama (v1.1b) Zhang et al. (2024) and chat variants of Llama2-7b and Llama2-13b et al. (2023). For all experiments, we aim to generate a set of five responses, i.e., $n=5$. The temperature value used in all corpus creation and generation experiments are fixed at 0.7. We do not use other decoding strategies. All experiments were conducted using a single A100 GPU.
@@ 5.3 Baselines
For framework evaluation, we utilize PC to generate a response set for each dialogue context in the test set. Subsequently, for PS, we use the fine-tuned ODRP model ($ODRP$) as well as the variant finetuned on hard negative samples ($ODRP$HN). Additionally, we introduce the following baseline response selection methods: 1) $rand$: Randomly selecting $r_{f}$ from $R_{n}$; 2) $cls$: Training a classifier (deberta-v2-large) from scratch with the curated preference dataset; 3) $pref$: Using the base OpenAssistant preference model without fine-tuning; 4) base LLM (either $TinyLlama$, $Llama2\text{-}7b$ or $Llama2\text{-}13b$): Generating a response by passing $D$ directly to the LLM i.e., standard LLM inference. Additionally, we leverage the zero-shot variant of the FS prompt (Figure 3) to generate a single response from both $Llama2\text{-}70b$ and $gpt\text{-}3.5\text{-}turbo$, allowing us to benchmark these against the responses produced by our framework when implemented with smaller LLMs.
@@ 5.4 Evaluation
Automatic Evaluation We evaluate the overall diversity and contextual coherence of the chosen responses by computing the inter-response Distinct-1,2 Li et al. (2016) and the UE-score Lee et al. (2022a) and UniEval-dialog coherence score Zhong et al. (2022) respectively. To evaluate the set of responses $R_{n}$ generated after MRG, we use several automatic metrics: inter-response semantic diversity ($d_{sem}$) and lexical diversity ($d_{lex}$) scores introduced in Section 3.2, as well as UE-score (UE), and UniEval-dialog coherence score (UniEval) to assess the quality of $R_{n}$. For inter-response diversity metrics, it should be highlighted that lower scores indicate greater lexical or semantic diversity.
Human Evaluation In our experiments, we also conduct a human evaluation to evaluate the efficacy of each PS approach. Similar to Smith et al. (2022); Sakaeda and Kawahara (2022), we engaged a group of five native english speaking participants for a comparative preference-based human evaluation. Each participant was presented a dialogue context along with a response generated by $ODRP$HN to compare against each of the other PS approaches ($base$, $rand$, $cls$, $ODRP$), as well as a response generated by $Llama2\text{-}70b$ and $gpt\text{-}3.5\text{-}turbo$, and told to select the agent they would rather converse with. Each participant was presented with 60 samples (30 from DailyDialog and 30 from EmpatheticDialogs) for each comparison. We report the Win, Tie and Loss percentage of each comparison.

### v2

@@ 1 Introduction
Prior research has primarily focused on modeling the o2m property using probabilistic learning frameworks, such as the Conditional Variational Auto-Encoder (CVAE) shen-etal-2017-conditional; zhao-etal-2017-learning, to enhance response diverity. These methods typically condition the response on both the dialogue context and a randomly sampled latent variable, capturing the variability in conversational responses and effectively modeling the o2m property. Other approaches include randomized architectures lee-etal-2022-randomized, Wasserstein Auto-encoders gu2018dialogwae, and Bayesian architectures lee2023empirical. These studies illustrate that while explicitly modeling the o2m property of OD significantly enhances response diversity, there is typically a trade-off with contextual coherence sun-etal-2021-generating; 9747458.
Recent advancements in Large Language Models (LLMs) have made it increasingly impractical to model the o2m property using probabilistic approaches, primarily due to the immense scale of modern LLMs zhao2023survey. These frameworks typically employ a pretrained LLM as the decoder, which is fine-tuned along with additional network components responsible for generating the latent distribution. This process becomes highly resource-intensive given the scale of these LLMs. Moreover, many state-of-the-art LLMs operate as black boxes with undisclosed parameters. Therefore, in the context of LLMs, adopting probabilistic frameworks for generating responses to model the o2m property has become largely impractical.
@@
To facilitate MRG, we curate o2mDial, a novel conversational dataset that explicitly captures the o2m property of OD. To create o2mDial, we leverage the DailyDialog corpus. First, we sample 500 dialogues (three to six turns) from the training set of the DailyDialog corpus. In this paper, for MRG, we fix $n=5$. In other words, we aim to generate a set of five lexically and semantically distinct, yet contextually coherent responses. Unlike prior datasets that feature multiple reference responses hedayatnia-etal-2022-systematic; sai-etal-2020-improving; gupta-etal-2019-investigating that rely on the same LLM to generate every reference response, we use five distinct LLMs to simulate five different agents, with each LLM generating one response. As far as possible, this ensures the semantic and lexical uniqueness of each response. Based on our resource constraints, we selected the following five LLMs: 1)gpt-3.5-turbo gpt3.5; 2)llama2-70b-chat touvron2023llama; 3)mixtral-8x22b jiang2024mixtral; 4) StableVicuna13b vicuna2023; 5) Flan-T5-xxl chung2022scaling. Additionally, to construct a separate test set for MRG evaluation, we sample another 100 dialogue samples from the test of the DailyDialog corpus. Similar to the training set, each turn consists of a dialogue context (three to six turns), and a set of five distinct and contextually coherent responses.
@@
# samples 500(train)/100(test)
@@ 3 Multi-Response Generation
It should be noted that even if all $n$ responses convey similar ideas, they can still be lexically unique by using different vocabularies or structures. Thus, a set of responses can exhibit high inter-response lexical diversity while maintaining low inter-response semantic diversity. For example, the statements "A heavy downpour is happening" and "There’s a strong rainfall occurring" are lexically unique due to their different words and structures. However, the semantic content—that it is raining heavily—remains the same, indicating high semantic similarity. Our goal is to generate responses with both high inter-response lexical diversity and high inter-response semantic diversity, while ensuring contextual coherence. Prior work in MRG involve utilizing different sampling strategies, which produce responses with low semantic diversity, and pre-specified dialogue acts, which are significantly more complex to implement sakaeda-kawahara-2022-generate.
@@ 3.1 Methods
Chain-of-thought (CoT) Prompt Chain-of-Thought (CoT) prompting wu-etal-2023-chain involves prompting the model to generate intermediate steps or explanations in addition to the final answer. In our case, we prompt the LLM to explain how each response differs from the other responses. We hypothesize that by prompting the model to identify the differences between generated responses, the model would be more inclined to generate lexically and semantically diverse responses. The prompt template is provided in Figure 4, located in the Appendix.
Prompt Chaining (PC) Prompt Chaining (PC) sun-etal-2024-prompt typically involves dividing a task into smaller subtasks and executing them sequentially using prompts, where the output of one prompt serves as the input for the next. In our approach, we use PC to guide the LLM in generating a set of $n$ unique responses one by one. The process begins with an initial prompt $\mathbf{P}_{0}$ that asks the LLM to generate a response to the dialogue context. Subsequent prompts ($\mathbf{P}_{1}\cdots\mathbf{P}_{n-1}$) instructs the LLM to generate contextually coherent responses that differ semantically and lexically from every response generated by the previous prompts, which are included in the current prompt as input. We hypothesize that by decomposing the task of MRG into $n$ smaller subtasks, the LLM can more effectively ensure both lexical and semantic uniqueness across the responses. However, it is important to note that PC requires multiple inferences from the LLM. Therefore, generating $n$ responses requires $n$ separate inferences, which could impact the feasibility and efficiency of this approach in the real world. The prompt template is provided in Figure 5, located in the Appendix.
@@
Instruction Tuning (IT) In addition, we also conduct IT via QLoRAdettmers2023qlora using the collected corpus. IT with QLoRA was performed using a batch size of 32, a learning rate of 2e-4, 4 epochs, a rank of 16, an alpha of 32, and a dropout of 0.05. The instruction used for IT is identical to the zero-shot variant (prompt consists of only the instruction without any demonstration examples) of the FS prompt (Figure 3).
@@ 3.2 Evaluation
Inter-response Semantic & Lexical Diversity In the context of open-domain dialogue, response diversity is typically measured via the Distinct metric, which is typically calculated by dividing the number of unique n-grams by the total number of n-grams. However, in our case, we aim to quantify the relative diversity of a set of $n$ responses. In other words, we would like to measure, on average, how different each response is from the other n1n-1 responses. Additionally, based on our definition, it would be ideal if semantic and lexical diversity can be evaluated separately. To this end, we define two separate metrics each accounting for either inter-response semantic or lexical diversity respectively: the inter-response semantic diversity score ($d_{sem}(R_{n})$) and inter-response lexical diversity score ($d_{lex}(R_{n})$).
@@
Contextual Coherence For our task, the overall contextual coherence of a set of responses can be attained by averaging the individual scores attained by each of the $n$ responses in $R_{n}$ (Algorithm 3). We employ two contextual coherence metrics: the Utterance Entailment (UE) score 9747458, and the UniEval-dialog coherence score zhong-etal-2022-towards. The UE score involves framing the task of contextual coherence evaluation as a Natural Language Inference (NLI) task. For each utterance in the dialogue context and the corresponding generated response, a NLI model assesses whether the response entails, contradicts, or is neutral with respect to the utterance. For each response $r_{i}$ from $R_{n}$, $UE(r_{i})=\frac{1}{m}\sum_{j\in m}NLI(r_{i},d_{j})$. The UE score of a set of $n$ responses is a continuous number between 0 and 1, where a greater value would indicate greater contextual coherence. The UniEval-dialog is a LLM-based approach which involve re-framing response evaluation as a boolean question and answer task. Essentially a LLM is finetuned and prompted to generate either ’Yes’ or ’No’ to the question: ’Is this a coherent response given the dialogue history?’. Hence, for the UniEval-dialogue coherence metric, each response is assigned a score of 1 if ’Yes’ is generated or 0 if ’No’ is generated.
However, evaluating the contextual coherence of OD dialogue responses remains a challenging problem and an active area of research due to the o2m property li-etal-2016-diversity. Hence, in our experiments, we conduct a human evaluation to further support our findings.
@@ 4 Preference-based Selection (PS)
PS involves selecting the final response $r_{f}$ from $R_{n}$ based on human preference. Unlike traditional open-domain dialogue criteria such as coherence, diversity, engagingness, naturalness, or fluency, human preference covers broader factors like helpfulness, harmlessness, and interestingness li-etal-2024-dissecting. We prioritize human preference for three key reasons. Firstly, MRG already ensures coherence and diversity within $R_{n}$. Secondly, existing metrics fall short in capturing the full complexity of human preferences, as they address only specific aspects of response quality jiang2024surveyhumanpreferencelearning. Thirdly, modern LLMs are largely capable of generating fluent, natural, and engaging responses zhao2023survey.
However, in addition to human preference, the contextual coherence of the response should still be considered during selection. Hence, for PS, we aim to design an Open-domain Dialogue Response Preference (ODRP) model that assigns a scalar score to each response in $R_{n}$ based on human preference. To achieve this, we leverage an open-source preference model from OpenAssistant deberta-v3-large-v2 based on deberta-v3-large commonly used for Reinforcement Learning with Human Feedback (RLHF) bai2022traininghelpfulharmlessassistant. Such models are typically trained on preference datasets derived from tasks such as summarization stienon2020learning and question answering nakano2021webgpt, or curated specifically to prevent harmful behavior bai2022traininghelpfulharmlessassistant. Hence, to fine-tune the preference model for open-domain dialogue, we construct a new preference dataset from the corpus described earlier.
Preference datasets consist of comparisons between two responses given the same prompt (dialogue context $D$ in our case). Extending o2mDial, we construct a preference dataset for fine-tuning by engaging annotators to label the preferred response $y_{c}$ and the rejected response $y_{r}$ (based on which they would prefer from a conversation partner) for every possible pair of responses from $R_{n}$, resulting in $n\choose 2$ pairs per set. As per ouyang2022training, we consider every pair from $R_{n}$ as a single batch. The preference model is then fine-tuned on the following contrastive loss function:
@@
We then proceed to fine-tune the preference model via QLoRA dettmers2023qlora for two epochs with AdamW (lr=2e-4). After MRG, the ODRP model assigns a score to each response in $R_{n}$, and $r_{f}$ is the response with the highest score:
@@
Additionally, we introduce a variant of the ODRP model finetuned on a subset of the corpus selected via hard negative sampling robinson2021contrastive. Specifically, we apply the base preference model to the dataset and deliberately extracted samples ( 50%) on which the base model performed the worst (assigned a similar score for both $y_{c}$ and $y_{r}$ or assigned a higher score to $y_{r}$). We finetuned this variant of the ODRP model for four epochs instead.
@@ 5.1 Corpora
For evaluation, we use two main datasets: DailyDialog li-etal-2017-dailydialog and EmpatheticDialogs rashkin-etal-2019-towards. DailyDialog features diverse, open-domain multi-turn conversations, while EmpatheticDialogs focuses on responses to emotionally grounded events. In our experiments, the dialogue agent’s task is to generate responses based solely on the context of the ongoing conversation. We do not use any additional information such as response labels (e.g., emotion, topic, or style) or speaker labels.
@@ 5.2 Implementation
We generate five responses per context ($n=5$) using TinyLlama (v1.1b) zhang2024tinyllamaopensourcesmalllanguage and chat variants of Llama2-7b and Llama2-13b touvron2023llama. For all experiments, we aim to generate a set of five responses, i.e., $n=5$. The temperature value used in all corpus creation and generation experiments are fixed at 0.7. We do not use other decoding strategies. All experiments were conducted using a single A100 GPU.
@@ 5.3 Baselines
For framework evaluation, we utilize PC to generate a response set for each dialogue context in the test set. Subsequently, for PS, we use the fine-tuned ODRP model ($ODRP$) as well as the variant finetuned on hard negative samples (ODRPODRPHN). Additionally, we introduce the following baseline response selection methods: 1) randrand: Randomly selecting $r_{f}$ from $R_{n}$; 2) clscls: Training a classifier (deberta-v2-large) from scratch with the curated preference dataset; 3) prefpref: Using the base OpenAssistant preference model without fine-tuning; 4) base LLM (either $TinyLlama$, $Llama2\text{-}7b$ or $Llama2\text{-}13b$): Generating a response by passing $D$ directly to the LLM i.e., standard LLM inference. Additionally, we leverage the zero-shot variant of the FS prompt (Figure 3) to generate a single response from both $Llama2\text{-}70b$ and $gpt\text{-}3.5\text{-}turbo$, allowing us to benchmark these against the responses produced by our framework when implemented with smaller LLMs.
@@ 5.4 Evaluation
Automatic Evaluation We evaluate the overall diversity and contextual coherence of the chosen responses by computing the inter-response Distinct-1,2 li-etal-2016-diversity and the UE-score 9747458 and UniEval-dialog coherence score zhong-etal-2022-towards respectively. To evaluate the set of responses $R_{n}$ generated after MRG, we use several automatic metrics: inter-response semantic diversity ($d_{sem}$) and lexical diversity ($d_{lex}$) scores introduced in Section 3.2, as well as UE-score (UE), and UniEval-dialog coherence score (UniEval) to assess the quality of $R_{n}$. For inter-response diversity metrics, it should be highlighted that lower scores indicate greater lexical or semantic diversity.
Human Evaluation In our experiments, we also conduct a human evaluation to evaluate the efficacy of each PS approach. Similar to smith-etal-2022-human; sakaeda-kawahara-2022-generate, we engaged a group of five native english speaking participants for a comparative preference-based human evaluation. Each participant was presented a dialogue context along with a response generated by ODRPODRPHN to compare against each of the other PS approaches (basebase, randrand, clscls, ODRPODRP), as well as a response generated by $Llama2\text{-}70b$ and $gpt\text{-}3.5\text{-}turbo$, and told to select the agent they would rather converse with. Each participant was presented with 60 samples (30 from DailyDialog and 30 from EmpatheticDialogs) for each comparison. We report the Win, Tie and Loss percentage of each comparison.